# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

### 11 Mar 2024 v2

@@ 2.1. Plücker weight vectors and weighted Grassmann orbifolds
Let $\Lambda^{k}(\mathbb{C}^{n})$ be the $k$-th exterior product of the complex $n$-dim space $\mathbb{C}^{n}$. Note that the standard ordered basis $\{e_{1},e_{2},\dots,e_{n}\}$ for $\mathbb{C}^{n}$ induces an ordered basis $\{e_{\lambda^{0}},e_{\lambda^{1}},\dots,e_{\lambda^{m}}\}$ of $\Lambda^{k}(\mathbb{C}^{n})$, where $e_{\lambda^{i}}=e_{\lambda^{i}_{1}}\wedge e_{\lambda^{i}_{2}}\wedge\dots\wedge e_{\lambda^{i}_{k}}$ for a schubert symbol $\lambda^{i}=(\lambda^{i}_{1},\lambda^{i}_{2},\dots,\lambda^{i}_{k})$. Then one can identify $\Lambda^{k}(\mathbb{C}^{n})$ with $\mathbb{C}^{m+1}$. Consider a subset of $\Lambda^{k}(\mathbb{C}^{n})$ defined by the following
@@
where $\lambda^{r_{j}}=(i_{1},...,i_{k-1},\ell_{j})$ and $\lambda^{s_{j}}=(\ell_{0},...,\widehat{\ell}_{j},...,\ell_{k})$ for any two ordered sequences $1\leq i_{1}<\dots<i_{k-1}\leq n$ and $1\leq\ell_{0}<\ell_{1}<\dots<\ell_{k}\leq n$, see [11, Theorem 3.4.11].
The relation described in (2.4) is known as a Plücker relation. The coordinates of $Pl(k,n)$ are known as Plücker coordinates.
@@
In (2.4), we can assume $2k\leq n$. If $2k>n$ consider $k^{\prime}=n-k$, then $2k^{\prime}=2n-2k<n$. Note that (${n\choose k}={n\choose k^{\prime}}$. Let $\lambda=(\lambda_{1},\dots,\lambda_{k})$ be a Schubert symbol for $k<n$. Then $\{1,2,\dots,n\}\setminus\{\lambda_{1},\dots,\lambda_{k}\}$ in a increasing order gives a Schubert symbol for $k^{\prime}<n$. This correspondence of the Schubert symbol gives a relation as in (2.4) in $Pl(k,n)$ induced from $Pl(k^{\prime},n)$.
Consider the ${\bf{b}}$-action of $\mathbb{C}^{*}$ on $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ defined in (1.1). Now the natural question is whether $Pl(k,n)$ is a $\mathbb{C}^{*}$-invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ under the ${\bf{b}}$-action defined in (1.1). This is not true in general for any arbitrary ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$. Note that $Pl(k,n)$ will be an invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ iff
@@
Consider two ordered sequences $1\leq i_{1}<\dots<i_{k-1}\leq n$ and $1\leq\ell_{0}<\ell_{1}<\dots<\ell_{k}\leq n$. Then we have $k+1$ pairs $(r_{0},s_{0}),\dots,(r_{k},s_{k})$. Consider two pair $(r_{i},s_{i})$, $(r_{j},s_{j})$ for $i\neq j$ and define $z=(z_{0},z_{1},\dots,z_{m})$ by $z_{r_{i}}=z_{s_{i}}=z_{r_{j}}=1$ and $z_{s_{j}}=(-1)^{i+j+1}$. Then $z\in Pl(k,n)$ and (2.5) reduced to $t^{b_{r_{j}}+b_{s_{j}}}=t^{b_{r_{i}}+b_{s_{i}}}$, for all $t\in\mathbb{C}^{*}$. This implies $b_{r_{i}}+b_{s_{i}}=b_{r_{j}}+b_{s_{j}}$. Thus both (2.4) and (2.5) satisfies together for any $t\in\mathbb{C}^{*}$ and for every $z\in Pl(k,n)$ iff $b_{r_{j}}+b_{s_{j}}$ is constant for every $j=0,1,\dots,k$. This motivates us to introduce the Plücker weight vector.
Note that a Plücker relation $\mathfrak{p}$ as in (2.4) uniquely corresponds to $(k+1)$ many pairs $\{(r_{0}^{\mathfrak{p}},s_{0}^{\mathfrak{p}}),(r_{1}^{\mathfrak{p}},s_{1}^{\mathfrak{p}}),\dots,(r_{k}^{\mathfrak{p}},s_{k}^{\mathfrak{p}})\}$. In the next definition and throughout the whole paper, we assume $k$ and $n$ are two positive integers with $k<n$ and $m+1={n\choose k}$.
@@
[…]
for all $j\in\{0,1,\dots,k\}$, where the $(k+1)$ many pairs $\{(r_{0}^{\mathfrak{p}},s_{0}^{\mathfrak{p}}),(r_{1}^{\mathfrak{p}},s_{1}^{\mathfrak{p}}),\dots,(r_{k}^{\mathfrak{p}},s_{k}^{\mathfrak{p}})\}$ are uniquely defined for every Plücker relation $\mathfrak{p}$.
@@
in the ordering as in (2.2). Consider $z=(z_{0},z_{1},\dots,z_{5})\in\mathbb{C}^{6}\setminus\{0\}$. Let
@@
Using Remark 2.2, it follows that $Pl(k,n)$ is $\mathbb{C}^{*}$-invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ under the ${\bf b}$-action iff ${\bf b}$ is a Plücker weight vector. This paves the way to define the following.
@@
Let $k$ and $n$ be two positive integers such that $k<n$ and $m+1={n\choose k}$. If ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ is a Plücker weight vector there is a ${\bf b}$-action of $\mathbb{C}^{*}$ on $Pl(k,n)$ as in (1.1). We denote the orbit space
@@
If ${\bf{b}}=(1,1,\dots,1)$ then ${\bf{b}}$ is trivially a Plücker weight vector. In this case the space $\mbox{Gr}_{\bf{b}}(k,n)$ is the Grassmann manifold $\mbox{Gr}(k,n)$. We denote the corresponding quotient map by
@@
Now we recall the definition of weighted Grassmannians from [1] and show that the space $\mbox{Gr}_{\bf b}(k,n)$ in Definition 2.6 is a generalization of $\mbox{WGr}(k,n)$ in [1, Definition 2.2].
Let $W:=(w_{1},w_{2},\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$. For $k<n$, recall $Pl(k,n)$ from (2.3). The algebraic torus (*$(\mathbb{C}^{*})^{n+1}$ acts on $Pl(k,n)$ by the following
(2.6) $(t_{1},t_{2},\dots,t_{n},t)\sum_{i=0}^{m}z_{i}e_{\lambda^{i}}=\sum_{i=0}^{m}t{\cdot}t_{\lambda^{i}}z_{i}e_{\lambda^{i}}$
Consider a subgroup $WD$ of (*$(\mathbb{C}^{*})^{n+1}$ defined by
@@
Then there is a restricted action of $WD$ on $Pl(k,n)$. Abe and Matsumura called the quotient $\frac{Pl(k,n)}{WD}$ a weighted Grassmannian and denote it by $\mbox{WGr}(k,n)$ in [1].
For all $i\in\{0,1,\dots,m\}$, let
@@
where $\lambda^{i}=(\lambda_{1}^{i},\lambda_{2}^{i},\dots,\lambda_{k}^{i})$ is the $i$-th Schubert symbol given in (2.2). Then $b_{i}\geq 1$ for all $i\in\{0,1,\dots,m\}$. Consider the weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$. From the Definition 2.3 it follows that ${\bf{b}}$ is a Plücker weight vector. Note that action of $WD$ on $Pl(k,n)$ is same as the ${\bf{b}}$-action of $\mathbb{C}^{*}$ on $Pl(k,n)$. Thus $\mbox{Gr}_{\bf{b}}(k,n)$ is same as the weighted Grassmannian $\mbox{WGr}(k,n)$, where the Plücker weight vector ${\bf{b}}=(b_{0},\dots,b_{m})$ is defined using (2.8).
Now, we check whether the converse is true. More explicitly, for a Plücker weight vector ${\bf{b}}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ whether there exist $W:=(w_{1},w_{2},\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$.
@@
For a given Plücker weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$, where $m+1={n\choose k}$ for $k<n$, there may not exist $W=(w_{1},w_{2}\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$. Let $n=4$ and $k=2$. Then $m={4\choose 2}-1=5$. In this case, we have $6$ Schubert symbols given by
Consider a weight vector ${\bf{b}}=(5,1,4,3,6,2)\in(\mathbb{Z}_{\geq 1})^{6}$. Then ${\bf{b}}$ is a Plücker weight vector follows from Example 2.5. If there exist $W=(w_{1},w_{2},w_{3},w_{4})\in(\mathbb{Z}_{\geq 0})^{4}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$ then $b_{1}=w_{1}+w_{3}+a=1$. Now $W\in(\mathbb{Z}_{\geq 0})^{4}$ implies that $b_{1}\geq a$. The only choice for $a$ is $1$. Then we have $w_{1}+w_{3}=0$. Also $w_{1}-w_{3}=b_{0}-b_{3}=2$. Which gives $w_{1}=1$, $w_{3}=-1$, $w_{2}=3$ and $w_{4}=2$. Note that here $W\notin(\mathbb{Z}_{\geq 0})^{4}$.
If we choose another Plücker weight vector ${\bf{b}}=(5,1,4,4,7,3)$, then only choice of $a=1$. Similar to the above calculation we get $W=(1/2,7/2,-1/2,5/2)$.
@@
Consider the linear equations defined by $w_{\lambda^{i}}=b_{i}$ for all $i=0,1,\dots,m$ with variables $w_{1},w_{2},\dots,w_{n}\text{ and }a$. Let us fix the variable $w_{1}$. For every $j$ satisfying $2\leq j\leq n$ we can choose two Schubert symbols $\lambda^{r_{j}}$ and $\lambda^{s_{j}}$ such that $(1,j)\lambda^{r_{j}}=\lambda^{s_{j}}$. Then we have linear equations of the form
@@
Note that if $(1,j)\lambda^{q_{j}}=\lambda^{t_{j}}$ for any two Schubert symbols $\lambda^{q_{j}}$ and $\lambda^{t_{j}}$ then $\lambda^{r_{j}}\cap\lambda^{t_{j}}=\lambda^{q_{j}}\cap\lambda^{s_{j}}$. Since ${\bf b}$ is a Plücker weight vector, using Remark 2.4, $b_{r_{j}}+b_{t_{j}}=b_{q_{j}}+b_{s_{j}}$. So, $b_{r_{j}}-b_{s_{j}}=b_{q_{j}}-b_{t_{j}}$. Thus (2.9) is well defined if ${\bf b}$ is a Plücker weight vector.
Now fix a Schubert symbol $\lambda^{i}=(2,3,\dots,k+1)$. We add the $k$ equations of (2.9) for $12\leq j\leq k+$ 1 with the equation $w_{\lambda^{i}}=b_{i}$ and we get
@@
For a given Plücker weight vector ${\bf{b}}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{n}$ there are infinitely many solutions of the form $W^{\prime}=(w_{1}-\alpha,\dots,w_{n}-\alpha)\in\mathbb{Z}^{n}$ and $a^{\prime}=a+k\alpha$ for any $\alpha\in\mathbb{Z}$ such that $b_{i}=w_{\lambda^{i}}$.
@@
2.2. $q$-CW complex structures of weighted Grassmann orbifolds
In this subsection, we discuss the $q$-CW complex structure of $\mbox{Gr}_{\bf{b}}(k,n)$ following [7, 1]. This $q$-CW complex structure induces a building sequence on $\mbox{Gr}_{\bf b}(k,n)$. Then, we explore the cofibrations that appear at each stage of the building sequence.
The CW complex structure of $\mbox{Gr}(k,n)$ given by
@@ 2.2. $q$-CW complex structures of weighted Grassmann orbifolds
${E}(\lambda^{i})=\{[(z_{0},z_{1},\dots,z_{m})]\in\mbox{Gr}(k,n)~{}|~{}z_{i}=1,z_{j}=0{\mbox{ for }}j>i\}.$
Every element of $E(\lambda^{i})$ has a presentation $[A]=[a_{1}\wedge a_{2}\wedge\dots\wedge a_{k}]$. If $\lambda^{i}=(\lambda^{i}_{1},\lambda^{i}_{2},\dots,\lambda^{i}_{k})$ is the $i$-th Schubert symbol then
(2.10) as:=($a_{s}:=(a_{s1},a_{s2},\dots,a_{s{\lambda^{i}_{s}}},\dots,0)\in\mathbb{C}^{n},$
such that ${a}_{s\lambda^{i}_{s}}=1$ and ${a}_{s\lambda^{i}_{j}}=0$ for $11\leq j\leq s-$ 1 and $1\leq s\leq k$, see [16].
Then $E(\lambda^{i})$ is an open cell of $\mbox{Gr}(k,n)$ of dimension $2d(\lambda^{i})$, where

### 09 Mar 2026 v3

@@ 2.1. Plücker weight vectors and weighted Grassmann orbifolds
Let $\Lambda^{k}(\mathbb{C}^{n})$ be the $k$-th exterior product of the $n$-dimensional complex vector space $\mathbb{C}^{n}$. Note that the standard ordered basis $\{e_{1},e_{2},\dots,e_{n}\}$ of $\mathbb{C}^{n}$ induces an ordered basis $\{e_{\lambda^{0}},e_{\lambda^{1}},\dots,e_{\lambda^{m}}\}$ of $\Lambda^{k}(\mathbb{C}^{n})$, where $e_{\lambda^{i}}=e_{\lambda^{i}_{1}}\wedge e_{\lambda^{i}_{2}}\wedge\dots\wedge e_{\lambda^{i}_{k}}$ for a Schubert symbol $\lambda^{i}=(\lambda^{i}_{1},\lambda^{i}_{2},\dots,\lambda^{i}_{k})$. Then one can identify $\Lambda^{k}(\mathbb{C}^{n})$ with $\mathbb{C}^{m+1}$. Consider a subset of $\Lambda^{k}(\mathbb{C}^{n})$ defined by the following
@@
where $\lambda^{r_{j}}=(i_{1},\ldots,i_{k-1},\ell_{j})$ and $\lambda^{s_{j}}=(\ell_{0},\ldots,\widehat{\ell}_{j},...,\ell_{k})$ for any two ordered sequences $1\leq i_{1}<\dots<i_{k-1}\leq n$ and $1\leq\ell_{0}<\ell_{1}<\dots<\ell_{k}\leq n$, see [13, Theorem 3.4.11].
The relation described in (2.4) is known as a Plücker relation. The coordinates of $Pl(k,n)$ are known as Plücker coordinates. Note that a Plücker relation $\mathfrak{p}$ as in (2.4) uniquely corresponds to $(k+1)$ many pairs $\{(r_{0}^{\mathfrak{p}},s_{0}^{\mathfrak{p}}),(r_{1}^{\mathfrak{p}},s_{1}^{\mathfrak{p}}),\dots,(r_{k}^{\mathfrak{p}},s_{k}^{\mathfrak{p}})\}$ of elements in $\{0,1,\dots,m\}$.
@@
In the definition of Plücker relation, we can assume $2k\leq n$. If $2k>n$ consider $k^{\prime}=n-k$, then $2k^{\prime}=2n-2k<n$. Note that ${n\choose k}={n\choose k^{\prime}}$. Let $\lambda=(\lambda_{1},\dots,\lambda_{k})$ be a Schubert symbol for $k<n$. Then $\{1,2,\dots,n\}\setminus\{\lambda_{1},\dots,\lambda_{k}\}$ in an increasing order gives a Schubert symbol for $k^{\prime}<n$. This correspondence of the Schubert symbol gives a relation as in (2.4) in $Pl(k,n)$ induced from $Pl(k^{\prime},n)$.
Consider the $\bf{b}$}-action of $\mathbb{C}^{*}$ on $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ defined in (1.1). Now the natural question is whether $Pl(k,n)$ is an invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ with respect to this $\bf{b}$}-action. Note that $Pl(k,n)$ will be an invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ iff
@@
This may not be true in general for any arbitrary ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$.
@@
Consider $(k+1)$ pairs $(r_{0},s_{0}),\dots,(r_{k},s_{k})$ corresponding to a Plücker relation. We fix two pairs $(r_{i},s_{i})$, $(r_{j},s_{j})$ for $i\neq j$ and consider $z=(z_{0},z_{1},\dots,z_{m})\in\mathbb{C}^{m+1}$ by $z_{r_{i}}=z_{s_{i}}=z_{r_{j}}=1$, $z_{s_{j}}=(-1)^{i+j+1}$ and the remaining components are zero. Then $z\in Pl(k,n)$ and (2.5) reduced to $t^{b_{r_{j}}+b_{s_{j}}}=t^{b_{r_{i}}+b_{s_{i}}}$, for all $t\in\mathbb{C}^{*}$. This implies $b_{r_{i}}+b_{s_{i}}=b_{r_{j}}+b_{s_{j}}$. Thus both (2.4) and (2.5) satisfies together for any $t\in\mathbb{C}^{*}$ and for every $z\in Pl(k,n)$ iff $b_{r_{j}}+b_{s_{j}}=b_{r_{i}}+b_{s_{i}}$ is constant for all $i,j$ in $\{0,1,\dots,k\}$. This motivates us to introduce the Plücker weight vector.
In the next definition and throughout the paper, we assume $k$ and $n$ are two positive integers with $k<n$ and $m+1={n\choose k}$.
@@
$b_{r_{j}^{\mathfrak{p}}}+b_{s_{j}^{\mathfrak{p}}}=b_{r_{i}^{\mathfrak{p}}}+b_{s_{i}^{\mathfrak{p}}}\text{ for all }i,j\text{ in }\{0,1,\dots,k\},$
where the $(k+1)$ many pairs $\{(r_{0}^{\mathfrak{p}},s_{0}^{\mathfrak{p}}),(r_{1}^{\mathfrak{p}},s_{1}^{\mathfrak{p}}),\dots,(r_{k}^{\mathfrak{p}},s_{k}^{\mathfrak{p}})\}$ are uniquely defined for every Plücker relation $\mathfrak{p}$.
@@
in the total ordering as in (2.2). Consider $z=(z_{0},z_{1},\dots,z_{5})\in\mathbb{C}^{6}\setminus\{0\}$. Let
@@
Using Remark 2.2, it follows that $Pl(k,n)$ is an invariant subset of $\mathbb{C}^{m+1}\setminus\{{\bf 0}\}$ under the $\bf b$}-action of $\mathbb{C}^{*}$ iff $\bf b$} is a Plücker weight vector. This paves the way for the definition of the following:
@@
Let ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ be a Plücker weight vector. We define a topological space $\mbox{Gr}_{\bf{b}}(k,n)$ as the orbit space of the $\bf b$}-action of $\mathbb{C}^{*}$ on $Pl(k,n)$ as in (1.1). Thus
@@
If ${\bf{b}}=(1,1,\dots,1)$ then $\bf{b}$} is a Plücker weight vector. In this case the space $\mbox{Gr}_{\bf{b}}(k,n)$ is the Grassmann manifold $\mbox{Gr}(k,n)$. We denote the corresponding quotient map by
@@
Now we show that the space $\mbox{Gr}_{\bf b}(k,n)$ in Definition 2.6 is a generalization of weighted Grassmannians $\mbox{WGr}(k,n)$ in [1, Definition 2.2]. Let $W:=(w_{1},w_{2},\dots,w_{n})\in(\mathbb{Z}_{\geq 0})^{n}$ and $a\in\mathbb{Z}_{\geq 1}$. For $k<n$, recall $Pl(k,n)$ from (2.3). The algebraic torus $(\mathbb{C}^{*})^{n+1}$ acts on $Pl(k,n)$ by the following
(2.6) $(t_{1},t_{2},\dots,t_{n},t)\sum_{i=0}^{m}z_{i}e_{\lambda^{i}}=\sum_{i=0}^{m}t{\cdot}t_{\lambda^{i}}z_{i}e_{\lambda^{i}}.$
Consider a subgroup $WD$ of $(\mathbb{C}^{*})^{n+1}$ defined by
@@
Then there is a restricted action of $WD$ on $Pl(k,n)$. Abe and Matsumura denoted the quotient $\frac{Pl(k,n)}{WD}$ by $\mbox{WGr}(k,n)$ and call it a weighted Grassmannian in [1]. For all $i\in\{0,1,\dots,m\}$, let
@@
where $\lambda^{i}=(\lambda_{1}^{i},\lambda_{2}^{i},\dots,\lambda_{k}^{i})$ is the $i$-th Schubert symbol given in (2.2). Then $b_{i}\geq 1$ for all $i\in\{0,1,\dots,m\}$. Consider the weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$. From the Definition 2.3, it follows that $\bf{b}$} is a Plücker weight vector. Note that action of $WD$ on $Pl(k,n)$ is same as the $\bf{b}$}-action of $\mathbb{C}^{*}$ on $Pl(k,n)$. Thus $\mbox{Gr}_{\bf{b}}(k,n)$ is same as the weighted Grassmannian $\mbox{WGr}(k,n)$, where the Plücker weight vector ${\bf{b}}=(b_{0},\dots,b_{m})$ is defined using (2.8).
@@
Let $n=4$ and $k=2$. Then $m={4\choose 2}-1=5$. Recall the $6$ Schubert symbols from Example 2.5. Consider a weight vector ${\bf{b}}=(5,1,4,3,6,2)\in(\mathbb{Z}_{\geq 1})^{6}$. Then $\bf{b}$} is a Plücker weight vector follows from Example 2.5. If there exist $W=(w_{1},w_{2},w_{3},w_{4})\in(\mathbb{Z}_{\geq 0})^{4}$ and $a\in\mathbb{Z}_{\geq 1}$ such that $b_{i}=w_{\lambda^{i}}$ then $b_{1}=w_{1}+w_{3}+a=1$. Now $W\in(\mathbb{Z}_{\geq 0})^{4}$ implies that $b_{1}\geq a$. The only choice for $a$ is $1$. Then we have $w_{1}+w_{3}=0$. Also $w_{1}-w_{3}=b_{0}-b_{3}=2$. Which gives $w_{1}=1$, $w_{3}=-1$, $w_{2}=3$ and $w_{4}=2$. Thus $W=(1,-1,3,2)\notin(\mathbb{Z}_{\geq 0})^{4}$.
If we choose another Plücker weight vector ${\bf{b}}=(5,1,4,4,7,3)$, then by a similar calculation as above, we get W=(1/2,7/2,1/2,5/2)W=(1/2,7/2,-1/2,5/2) and $a=1$. Moreover, $\bf b$} also corresponds to W=(0,3,1,2)W=(0,3,-1,2) and $a=2$.
@@
Consider the linear equations defined by $w_{\lambda^{i}}=b_{i}$ for all $i=0,1,\dots,m$ with variables $w_{1},w_{2},\dots,w_{n}\text{ and }a$. Let us fix the variable $w_{1}$. For every $j$ satisfying $2\leq j\leq n$ we can choose two Schubert symbols $\lambda^{r_{j}}$ and $\lambda^{s_{j}}$ such that $(1,j)\in\text{inv}(\lambda^{r_{j}})$ and $(1,j)\lambda^{r_{j}}=\lambda^{s_{j}}$. Then we have linear equations of the form
@@
Note that if there exists two Schubert symbols $\lambda^{q_{j}}$ and $\lambda^{t_{j}}$ such that $(1,j)\in\text{inv}(\lambda^{q_{j}})$ and $(1,j)\lambda^{q_{j}}=\lambda^{t_{j}}$ for then $\lambda^{r_{j}}\cup\lambda^{t_{j}}=\lambda^{q_{j}}\cup\lambda^{s_{j}}$. Since $\bf b$} is a Plücker weight vector, using Remark 2.4, $b_{r_{j}}+b_{t_{j}}=b_{q_{j}}+b_{s_{j}}$. So, $b_{r_{j}}-b_{s_{j}}=b_{q_{j}}-b_{t_{j}}$. Thus (2.9) is well defined if $\bf b$} is a Plücker weight vector.
Now fix a Schubert symbol $\lambda^{i}=(2,3,\dots,k+1)$. We add the $k$ equations of (2.9) for $2\leq j\leq k+1$ with the equation $w_{\lambda^{i}}=b_{i}$ and we get
@@
The Plücker weight vector corresponding to $W=(w_{1},\dots,w_{n})\in\mathbb{Z}^{n}$ and $a\in\mathbb{Z}$ is same as the Plücker weight vector corresponding to $W_{\alpha}=(w_{1}-\alpha,\dots,w_{n}-\alpha)\in\mathbb{Z}^{n}$ and $a_{\alpha}=a+k\alpha$ for any $\alpha\in\mathbb{Z}$.
@@
$2.$ $q$-CW complex structures of weighted Grassmann orbifolds
In this subsection, we discuss a $q$-CW complex structure of $\mbox{Gr}_{\bf{b}}(k,n)$. This $q$-CW complex structure induces a building sequence on $\mbox{Gr}_{\bf b}(k,n)$ and we explore the cofibrations that appear at each stage of the building sequence. A CW complex structure of $\mbox{Gr}(k,n)$ given by
@@
(2.10) ${E}(\lambda^{i})=\{[(z_{0},z_{1},\dots,z_{m})]\in\mbox{Gr}(k,n)~|~z_{i}=1,z_{j}=0{\mbox{ for }}j<i\}.$
Let $\lambda^{i}=(\lambda^{i}_{1},\lambda^{i}_{2},\dots,\lambda^{i}_{k})$ denote the $i$-th Schubert symbol. Then every element of $E(\lambda^{i})$ has a presentation $A]=[a_{1}\wedge a_{2}\wedge\dots\wedge a_{k}]$, where
(2.11) $a_{s}:=(a_{s1},a_{s2},\dots,a_{sn})\in\mathbb{C}^{n},$
such that ${a}_{s\lambda^{i}_{s}}=1$ and ${a}_{sj}=0$ for $j<\lambda^{i}_{s}$ and $j\in\{\lambda^{i}_{s+1},\dots,\lambda^{i}_{k}\}$, see [12, Section 1.1].
Then $E(\lambda^{i})$ is an open cell of $\mbox{Gr}(k,n)$ of codimension $2d(\lambda^{i})$, where
@@
Consider $\widetilde{E}(\lambda^{i}):=\pi^{-1}(E(\lambda^{i}))\subset Pl(k,n)$. Then $Pl(k,n)=\bigsqcup_{i=0}^{m}\widetilde{E}(\lambda^{i})$. Moreover,
$\widetilde{E}(\lambda^{i})=\{[(z_{0},z_{1},\dots,z_{m})]\in Pl(k,n)~|~z_{i}\neq 0,z_{j}=0{\mbox{ for }}j<i\}.$
Then $\widetilde{E}(\lambda^{i})$ is homeomorphic to ${E}(\lambda^{i})\times\mathbb{C}^{*}$ via the map
The inverse map of $f$ is defined by
The $\bf b$-action of $\mathbb{C}^{*}$ on $Pl(k,n)$ restricts a $\bf b$-action of $\mathbb{C}^{*}$ on $\widetilde{E}(\lambda^{i})$. Define a $\bf b$-action of $\mathbb{C}^{*}$ on $E(\lambda^{i})$ by
[…]
This gives rise to a $\bf{b}$}-action of $\mathbb{C}^{*}$ on ${E}(\lambda^{i})\times\mathbb{C}^{*}$ defined by
$t([z],s)=([tz],t^{b_{i}}s).$
Then $f$ is $\mathbb{C}^{*}$-equivariant with respect to the above $\bf{b}$}-actions. Let $G(b_{i})$ be the finite subgroup of $\mathbb{C}^{*}$ defined by
$G(b_{i})=\{t\in\mathbb{C}^{*}~|~t^{b_{i}}=1\}.$
Thus using [7, Lemma 2.6]
$\pi_{\bf{b}}(\widetilde{E}(\lambda^{i}))\cong\frac{{E}(\lambda^{i})\times\mathbb{C}^{*}}{{\bf b}\text{-action}}\cong\frac{{E}(\lambda^{i})}{G(b_{i})}.$
Therefore a $q$-cell decomposition of $\mbox{Gr}_{\bf b}(k,n)$ is given by $\mbox{Gr}_{\bf{b}}(k,n)=\sqcup_{i=0}^{m}\frac{E(\lambda^{i})}{G(b_{i})}.$
2.3. Lens complex and induced cofibration
Next, we recall lens complex from [15]. Let $(c_{0},c_{1},\dots,c_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$. For each $i\in\{0,1,\dots,m\}$ consider a unit sphere
$S^{2m-1}_{i}=\{(z_{0},\dots,z_{i-1},0,z_{i}\dots,z_{m})\in\mathbb{C}^{m+1}\setminus\{{\bf 0}\}~|~\sum_{j\neq i;~j=0}^{m}|z_{j}|^{2}=1\}.$
Definition 2.11.
[15] Consider an action of $G(c_{i})$ on $S^{2m-1}_{i}$ defined by
(2.13) $t(z_{0},z_{1},\dots,z_{i-1},0,z_{i+1},\dots,z_{m})=(t^{c_{0}}z_{0},t^{c_{1}}z_{1},\dots,t^{c_{i-1}}z_{i-1},0,t^{c_{i+1}}z_{i+1},\dots,t^{c_{m}}z_{m}),$
where $t\in G(c_{i})$. The quotient space
$L^{\prime}(c_{i};(c_{0},c_{1},\dots,c_{i-1},c_{i+1},\dots,c_{m})):=\dfrac{S^{2m-1}_{i}}{G(c_{i})}$
is called the lens complex. If $\{c_{j}:0\leq j\leq m,j\neq i\}$ are prime to $c_{i}$ then $L^{\prime}(c_{i};(c_{0},\dots,c_{i-1},$ $c_{i+1},\dots,c_{m}))$ is known as usual generalized lens space.
Next, we show that the lens complex arises naturally as a cofiber in the building sequence induced from the $q$-CW complex structure of $\mbox{Gr}_{\bf b}(k,n)$. For each $i\in\{0,1,\dots,m\}$, define $X^{m-i}:=\bigsqcup_{j\geq i}\frac{E(\lambda^{j})}{G(b_{j})}.$ The building sequence corresponding to the $q$-CW complex structure of $\mbox{Gr}_{\bf b}(k,n)$ is